OpenStudy (anonymous):

5x+3y=34 and -5x+2y=-19 can u show the working out for when ur getting y lease

OpenStudy (anonymous):

5x+3y=34 (+) -5x+2y=-19 => 5y=15 => y=3

OpenStudy (anonymous):

x=(34-3*3)/5= 25/5 = 5 so x=5;y=3